OpenStudy (anonymous):

A certain radioactive element decays over time according to the equation y=A(1/100)^t/300 where A= the number of grams present initially and t=time in years. If 1000 grams were present initially, how many grams will remain after 900 years?

OpenStudy (anonymous):

please help(:

OpenStudy (anonymous):

you are aware these equations are type first order reactions, right?

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

Chemical kinetics. NVM, we'll simplify the equation before solving it. Take log on both sides.

OpenStudy (anonymous):

how would I set it up exactly?

OpenStudy (anonymous):

that gives us lig (y) = (t/300)*log(1/100) + log A

OpenStudy (anonymous):

which gives log y= log A- 2(t/300)

OpenStudy (anonymous):

differenciating w.r.t. T gives us the rate of decay.

OpenStudy (anonymous):

wouldnt't I just plug in the values to the equation? like a and t?

OpenStudy (anonymous):

the value of y in the eqn gives ammount remaining and the ammount decayed would be by subtracting from the inintial..

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

sorry, I thought you needed to take the rate of reaction. =P Just put the values there and solve. That gives you the value required.

OpenStudy (anonymous):

1000 (1/100)^900/300 =1000*(1/100)^3 =0.001 is what is left.

OpenStudy (anonymous):

thats what I got and its not one of the answer choices... here ill post the choices

OpenStudy (anonymous):

\[y=A\left(1\over100\right)^{t\over300}\] plug in A = 1000 and t = 900

OpenStudy (anonymous):

these are the choices A) 500 grams B) 250 grams C)125 grams D) 62.5 grams

OpenStudy (anonymous):

@sparky16 is that how the equation looks? if not, use the equation editor below to enter it clearly.

OpenStudy (anonymous):

oh wait I typed it wrong...lemme fix it in the equation editor

OpenStudy (anonymous):

\[y=A(\frac{ 1 }{ 2 })^{\frac{ t }{ 100 }}\]

OpenStudy (anonymous):

100(1/2)9 which is 25/32

OpenStudy (anonymous):

pardon 125/16.

OpenStudy (anonymous):

@sparky16 I still think you've got it wrong

OpenStudy (anonymous):

is it perhaps like this: \[y=A\left(1\over2\right)^{t\over300}\]

OpenStudy (anonymous):

1.953125

OpenStudy (anonymous):

\[y=A\left(1\over2\right)^{\Large\color{red}{t\over300}}\]

OpenStudy (anonymous):

yes. it is.. my bad.

OpenStudy (anonymous):

ok, now plug in the numbers for "A" and "t"

OpenStudy (anonymous):

125?

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

be careful when you note down the equation.

OpenStudy (anonymous):

@NeetziD do not get disappointed. you would have got it right, if the equation was not mis-leading.

OpenStudy (anonymous):

No, I'm not. Thanks, by the way. I didn't read the requirement, and since chemistry is too high... been taking it in that direction. =P